# A Computationally Efficient Maximum A Posteriori Sequence Estimation via Stein Variational Inference

### v3

@@ Abstract
State estimation in robotic systems presents significant challenges, particularly due to the prevalence of multimodal posterior distributions in real-world scenarios. One effective strategy for handling such complexity is to compute maximum a posteriori (MAP) sequences over a discretized or sampled state space, which enables a concise representation of the most likely state trajectory. However, this approach often incurs substantial computational costs, especially in high-dimensional settings. In this article, we propose a novel MAP sequence estimation method, Stein-MAP-Seq, which effectively addresses multimodality while substantially reducing computational and memory overhead. Our key contribution is a sequential variational inference framework that captures temporal dependencies in dynamical system models and integrates Stein variational gradient descent (SVGD) into a Viterbi-style dynamic programming algorithm, enabling computationally efficient MAP sequence estimation. Stein-MAP-Seq achieves a computational complexity of $\mathcal{O}(M^{2})$, where $M$ is the number of particles, in contrast to the $\mathcal{O}(N^{2})$ complexity of conventional MAP sequence estimators, with $N\gg M$. Furthermore, the method inherits SVGD’s parallelism, enabling efficient computation for real-time deployment on GPU-equipped autonomous systems. We validate the proposed method in various multimodal scenarios, including those arising from nonlinear dynamics with ambiguous observations, unknown data associations, and temporary unobservability, demonstrating substantial improvements in estimation accuracy and robustness to multimodality over existing approaches.
Index Terms:
Variational inference, maximum a posteriori (MAP) sequence estimation, Viterbi algorithm, dynamic programming, Stein variational gradient descent.
@@ I Introduction
State estimation in real-world robotic applications requires a reliable and efficient estimator. This paper focuses on the design of computationally efficient and accurate Maximum A Posteriori (MAP) sequence estimators for robotic systems. MAP sequence (MAP-Seq) estimation, due to its ability to handle complex, non-Gaussian distributions and leverage temporal dependencies in dynamical systems, is often sought after in estimation scenarios involving multimodal posterior distributions. These distributions frequently arise in challenging robotics tasks due to nonlinear system-environment interactions, ambiguous observations, uncertain data associations, and multimodal sensor fusion.
@@
MAP-Seq estimators are an extension of MAP estimators, which seek the most probable state in a given distribution. MAP estimators are often preferred over Minimum Mean Square Error (MMSE) estimators for multimodal distributions, as MMSE estimators can produce estimates that lie between modes, potentially corresponding to low-probability regions and leading to erroneous state estimates. In contrast, MAP estimation seeks the most probable state, effectively capturing one of the dominant modes in multimodal distributions. The MAP-Seq estimator extends this concept to find the most likely trajectory of states over a given time horizon, conditioned on the entire sequence of observations. This approach offers more robust estimates, particularly in scenarios with temporary unobservability or when the system’s dynamics provide strong temporal constraints, by leveraging information across time to resolve ambiguities that may be present in point estimates. This framework has found applications across diverse domains, including image processing [1], natural language processing [2], and robotics for tasks such as high-level decision-making [3] and simultaneous localization and mapping (SLAM) or tracking [4, 5].
@@
While MAP-Seq estimation offers significant advantages, computing the optimal trajectory can be computationally expensive, especially for complex, high-dimensional systems. One common approach uses smoothing techniques, such as iterated Extended Kalman Smoothing (iEKS) and factor graph optimization [6], which iteratively estimate the trajectory under Gaussian assumptions and local linearization. However, these methods often struggle with multimodal distributions, as they are prone to converging to locally optimal trajectories [5]. In contrast, approaches based on finite (discrete or sampled) state spaces, such as the Viterbi algorithm, compute the globally optimal trajectory via dynamic programming for MAP-Seq estimation [7], thereby avoiding local optima. This makes them particularly well-suited for estimation problems involving ambiguity, non-Gaussianity, or multimodality.
@@
Despite this advantage, its computational complexity, which scales quadratically with the number of discrete states ($\mathcal{O}(N^{2})$ for $N$ states), poses a significant bottleneck for real-time applications, especially when dealing with systems requiring fine-grained discretization of the state space for accurate representation. This limitation often restricts Viterbi algorithm’s applicability to coarse-grid discretizations or necessitates the use of high-performance computing resources and substantial memory allocation in large-scale scenarios.
@@
Particle Filter-based approaches for MAP-Seq estimation have also been investigated [16, 17, 18, 19, 20]. While these methods have shown promise in handling complex multimodal distributions, they still rely on sequential Monte Carlo methods, which require many particles and suffer from resampling overhead and random sampling variability.
@@
In recent years, variational inference has emerged as a promising alternative for state estimation in complex systems. Variational methods aim to approximate intractable posterior distributions with simpler, tractable forms by framing inference as an optimization problem [21]. These techniques have gained attention in robotics and autonomous systems due to their ability to balance computational efficiency with the flexibility to capture complex, multimodal distributions [22, 23]. Notably, advances in variational inference have led to the development of methods that can handle non-Gaussian posteriors and temporal dependencies, making them particularly suitable for sequential estimation tasks [24, 25, 26]. The application of variational inference to state estimation problems has shown promise in various domains, including visual-inertial odometry [27], SLAM [28], and multi-target tracking [29].
@@
Statement of contribution: This paper presents Stein-MAP-Seq, a novel particle-based approach for MAP-Seq estimation in dynamical systems that addresses the challenges of multimodal state estimation while maintaining computational efficiency. Our method leverages recent advances in variational inference, specifically Stein Variational Gradient Descent (SVGD) [30], a powerful non-parametric variational inference method. SVGD works by deterministically evolving a set of particles that serve as samples from the target distribution, minimizing the Kullback-Leibler divergence between the particle-based empirical distribution and the true posterior. Unlike traditional Monte Carlo methods that generate independent samples through stochastic processes, SVGD produces a deterministic set of interacting particles that collectively capture the structure of complex, potentially multimodal distributions.
@@
The key innovation of Stein-MAP-Seq lies in leveraging these SVGD-generated particles as representative samples to transform the intractable continuous MAP optimization into an efficient discrete combinatorial problem. Stein-MAP-Seq embeds SVGD into a Viterbi-style dynamic programming algorithm within a sequential variational inference framework that explicitly models the temporal dependencies inherent in dynamical systems. This novel combination uniquely leverages SVGD’s mode-seeking capability to deterministically generate a compact set of samples that concentrate around the modes of the posterior distribution, enabling efficient MAP sequence inference via dynamic programming over this discrete sample space using log probability-based scoring and backtracking. This approach effectively captures the complexity of multimodal posteriors while achieving a computational complexity of $\mathcal{O}(M^{2})$, where $M$ is the number of particles, a significant improvement over the $\mathcal{O}(N^{2})$ cost of the conventional MAP-Seq estimators with $N\gg M$. Furthermore, since the particle updates in SVGD can be computed independently for each sample, SVGD’s inherent parallel nature makes Stein-MAP-Seq particularly well-suited for real-time deployment on autonomous systems with parallel computing architectures [31, 32, 33, 34, 35].
@@
Our approach fundamentally differs from traditional sampling-based methods by using SVGD’s deterministic particle evolution to generate high-quality samples that are then used as discrete states in a Viterbi-style dynamic programming algorithm. We rigorously evaluate the proposed method on challenging scenarios, including nonlinear dynamics with ambiguous measurements, pose estimation under unknown data associations, and real-world range-only (wireless) localization under temporary unobservability, demonstrating substantial improvements in estimation accuracy and robustness to multimodality over state-of-the-art baselines.
@@
The rest of this paper is organized as follows: Section II formally defines the problem and presents our assumptions. Section III provides a brief overview of SVGD. Section IV details our proposed Stein-MAP-Seq algorithm. Section V presents estimation results across diverse scenarios and compares them with existing methods. Finally, Section VI presents the conclusion.
@@ II Problem Definition
Solving (2) is computationally intractable, particularly when $p(x_{0:T}|z_{1:T})$ exhibits multimodality and/or lacks a closed-form expression, making exact analytical computation impossible. To address this challenge, we propose a sequential variational inference to obtain an approximate MAP sequence solution. The details of the proposed MAP formulation are presented in Section IV. Before presenting those details, the next section provides background on the Stein Variational Gradient Descent method, which is one of the key components of our approach.
@@ III A Brief Overview of Stein Variational Gradient Descent method
Variational inference (VI) approximates a target distribution $p(x)$ using a simpler distribution $q^{\star}(x)$ found in a predefined set $\mathcal{Q}$ by minimizing the Kullback–Leibler (KL) divergence [36], i.e.,
@@
where the choice of the model space $\mathcal{Q}$ is critical for balancing accuracy and computational efficiency; however, achieving this balance remains challenging. [^1] KL divergence is defined as $\text{KL}[q(x)\parallel p(x)]=\int q(x)\log\frac{q(x)}{p(x)}dx$.
@@
The update $x_{i}^{(l+1)}=x_{i}^{(l)}+\epsilon\hat{\phi}^{\star}(x_{i}^{(l)})$ iteratively pushes particles toward $p(x)$, enabling expectation approximations via the empirical mean. Algorithm 1 summarizes the Stein Variational Gradient Descent algorithm.
Input: The score function $\nabla_{x}\log p(x)$.
Goal: A set of particles $\{x_{i}\}_{i=1}^{N_{s}}$ that approximates $p(x)$.
For iteration l do
where $\hat{\phi}^{\star}(x_{i}^{(l)})$ is given in (III).
Algorithm 1 Stein Variational Gradient Descent
The samples generated by SVGD, denoted as $\{x_{i}\}_{i=1}^{N_{s}}$, serve a dual purpose in variational inference. First, they provide an empirical approximation of the optimal distribution $q^{\star}(x)$ in (3) as $q^{\star}(x)=\frac{1}{N_{s}}\sum\nolimits_{i=1}^{N_{s}}\delta(x-x^{i})$. Second, and equally important, these particles are commonly interpreted as representative samples from the target distribution $p(x)$ itself. Unlike traditional sampling methods such as MC or MCMC, which generate independent samples through stochastic processes, SVGD produces a deterministic set of interacting particles that collectively capture the structure of complex, potentially multimodal distributions. This deterministic particle-based representation offers computational advantages for downstream tasks, including efficient parallel computation and natural handling of multimodal posterior structures. Consequently, SVGD has emerged as a powerful sampling tool for approximate inference in high-dimensional and multimodal scenarios where traditional sampling methods face significant challenges.
Conventional SVGD addresses stationary problems, whereas we extend it to sequential inference in a state estimation framework, as discussed in the next section.
@@ IV Stein Maximum A Posteriori Sequence Estimation
In this section, we aim to develop a MAP sequence estimator for the dynamical system (1) by integrating a variational inference framework with dynamic programming. Computing the MAP trajectory $x_{0:T}^{\mathrm{MAP}}$ requires constructing an optimal distribution $q^{\star}(x_{0:T})$ whose support must be concentrated on the (unknown) global maxima of the true posterior $p(x_{0:T}|z_{1:T})$, and subsequently searching for these modes, making this task highly challenging. Formally, the problem can be stated as:
@@
Solving these optimization problems exactly presents significant challenges. Particularly, when the dynamical system (1) is nonlinear, the resulting true posterior $p(x_{0:T}|z_{1:T})$ becomes non-Gaussian and often multimodal, making its representation in closed-form impractical and exact analytical computation impossible. Moreover, the choice of $q(x_{0:T})$ significantly impacts the performance of the estimation [16]. In the second optimization problem, finding the mode of $q^{\star}(x_{0:T})$ is not tractable, since $q^{\star}(x_{0:T})$ typically lives on a high-dimensional continuous support.
To address these limitations, we propose Stein-MAP-Seq, a MAP-sequence estimator for dynamical systems (1) that preserves temporal dependencies via a sequential variational-inference framework (Lemma IV.2). Stein-MAP-Seq integrates SVGD into a Viterbi-style dynamic programming framework, where SVGD’s gradient flow provides a compact discrete state support, enabling efficient global MAP trajectory recovery via the Viterbi algorithm’s forward recursion and backtracking.
Our approach addresses the intractability of the continuous optimization in (5) through a two-stage strategy: first, we use SVGD to generate particle-based approximations of the sequential posterior distributions, transforming the continuous state space into a finite discrete support; second, we apply dynamic programming over this discrete particle support to efficiently find the globally optimal MAP trajectory. This particle-based discretization fundamentally changes the nature of the optimization from an intractable continuous search to an efficient discrete combinatorial problem.
The remainder of this section details the theoretical foundations and algorithmic steps of Stein-MAP-Seq, demonstrating how the two-stage combination of SVGD and Viterbi-style dynamic programming yields a robust and efficient MAP-sequence estimator.

### v4

@@ Abstract
State estimation in robotic systems presents significant challenges, particularly due to the prevalence of multimodal posterior distributions in real-world scenarios. One effective strategy for handling such complexity is to compute maximum a posteriori (MAP) sequences over a discretized or sampled state space, which enables a concise representation of the most likely state trajectory. However, this approach often incurs substantial computational costs, especially in high-dimensional settings. In this article, we propose a novel MAP sequence estimation method, Stein-MAP-Seq, which effectively addresses multimodality while substantially reducing computational and memory overhead. Our key contribution is a sequential variational inference framework that captures temporal dependencies in dynamical system models and integrates Stein variational gradient descent (SVGD) into a Viterbi-style dynamic programming algorithm, enabling computationally efficient MAP sequence estimation. This integration allows the method to focus computational effort on MAP-consistent modes rather than exhaustively exploring the entire state space. Stein-MAP-Seq inherits the parallelism and mode-seeking behavior of SVGD, allowing particle updates to be efficiently executed on parallel hardware and significantly reducing the number of trajectory candidates required for MAP-sequence recursion compared to conventional methods that rely on hundreds to thousands of particles. We validate the proposed approach on a range of highly multimodal scenarios, including nonlinear dynamics with ambiguous observations, unknown data association with outliers, range-only localization under temporary unobservability, and high-dimensional robotic manipulators. Experimental results demonstrate substantial improvements in estimation accuracy and robustness to multimodality over existing estimation methods.
@@ I Introduction
State estimation in robotic systems presents fundamental challenges, particularly when posterior distributions exhibit multimodality—a phenomenon that frequently arises in real-world applications due to nonlinear system-environment interactions, ambiguous observations, uncertain data associations, and multimodal sensor fusion. To address such complexity, Maximum A Posteriori (MAP) sequence estimators offer a principled framework that not only handles non-Gaussian distributions but also leverages temporal dependencies inherent in dynamical systems to produce trajectory-consistent estimates. This paper focuses on the design of a computationally efficient and accurate MAP sequence (MAP-Seq) estimator that effectively manages multimodal posterior distributions while maintaining practical computational requirements for robotic applications.
@@
MAP-Seq estimators are an extension of MAP estimators, which seek the most probable state in a given distribution. For multimodal distributions, MAP estimators are often preferred over Minimum Mean Square Error (MMSE) estimators, as MMSE estimators can produce estimates that lie between modes—potentially corresponding to low-probability regions that lead to erroneous state estimates. In contrast, MAP estimation seeks the most probable state, effectively capturing one of the dominant modes. MAP-Seq estimation extends this concept to find the most likely trajectory of states over a given time horizon, conditioned on the entire sequence of observations. By leveraging information across time to resolve ambiguities that may be present in point estimates, this approach offers more robust estimates, particularly in scenarios with temporary unobservability or when the system’s dynamics provide strong temporal constraints. This framework has found applications across diverse domains, including image processing [1], natural language processing [2], and robotics for tasks such as high-level decision-making [3] and simultaneous localization and mapping (SLAM) or tracking [4, 5].
@@
While MAP-Seq estimation offers significant advantages, computing the optimal trajectory can be computationally expensive, especially for complex, high-dimensional systems. Common approaches use smoothing techniques, such as iterated Extended Kalman Smoothing (iEKS) and factor graph optimization [6], which iteratively estimate the trajectory under Gaussian assumptions and local linearization. However, these methods often struggle with multimodal distributions, as they are prone to converging to locally optimal trajectories with poor initialization [5]. In contrast, approaches based on finite (discrete or sampled) state spaces, such as the Viterbi algorithm, compute the globally optimal trajectory via dynamic programming for MAP-Seq estimation [7], thereby avoiding local optima. This makes them particularly well-suited for estimation problems involving ambiguity, non-Gaussianity, or multimodality.
@@
Despite this global optimality advantage, the Viterbi algorithm’s computational complexity, which scales quadratically with the number of discrete states ($\mathcal{O}(N^{2})$ for $N$ states), poses a significant bottleneck for real-time applications, especially when dealing with systems requiring fine-grained discretization of the state space for accurate representation. This limitation often restricts the Viterbi algorithm’s applicability to coarse-grid discretizations or necessitates the use of high-performance computing resources and substantial memory allocation in large-scale scenarios.
@@
As an alternative to discrete-state Viterbi approaches, Particle Filter-based methods for MAP-Seq estimation have also been investigated [16, 17, 18, 19, 20]. While these methods have shown promise in handling complex multimodal distributions, they still rely on sequential Monte Carlo sampling, which requires large particle sets and suffers from resampling overhead and random sampling variability.
Recent advances in multimodal state estimation have explored deep learning-based paradigms, including amortized variational inference for nonlinear state-space models that learn neural networks to approximate filtering distributions [21], neural particle filtering approaches that incorporate end-to-end learned models [22], and diffusion-based generative models for state estimation that perform posterior inference via iterative denoising processes [23]. These methods exhibit strong representational capacity for complex and multimodal posteriors; however, they typically rely on offline training and implicit inference models, for which posterior densities are not explicitly available, making principled MAP-Seq estimation difficult.
@@
The computational and representational challenges inherent in MAP-Seq estimation—including the limitations of discrete-state methods, sampling variability in particle filters, and implicit posteriors in learning-based approaches—motivate the need for alternative frameworks. Variational inference offers a promising avenue to address these challenges. Variational methods aim to approximate intractable posterior distributions with simpler, tractable forms by framing inference as an optimization problem [24]. Unlike the implicit representations used in deep learning approaches, variational inference provides explicit posterior approximations that enable principled MAP-Seq computation. These techniques have gained attention in robotics and autonomous systems due to their ability to balance computational efficiency with the flexibility to capture complex, multimodal distributions [25, 26]. Notably, advances in variational inference have led to the development of methods that can handle non-Gaussian posteriors and temporal dependencies, making them particularly suitable for sequential estimation tasks [27, 28, 29]. The application of variational inference to state estimation problems has shown promise in various domains, including visual-inertial odometry [30], SLAM [31], and multi-target tracking [32].
@@
Statement of contribution: This paper presents Stein-MAP-Seq, a novel particle-based approach for MAP-Seq estimation in dynamical systems that addresses the challenges of multimodal state estimation while maintaining computational efficiency. Our method leverages Stein Variational Gradient Descent (SVGD) [33], a powerful non-parametric variational inference technique that deterministically evolves a set of particles to serve as samples from the target distribution, minimizing the Kullback-Leibler divergence between the particle-based empirical distribution and the true posterior. Unlike traditional Monte Carlo methods that rely on stochastic sampling and/or resampling procedures, SVGD deterministically transports particles toward high-probability regions using functional gradients that collectively capture the structure of complex, potentially multimodal distributions.
@@
The key innovation of Stein-MAP-Seq lies in leveraging SVGD-generated particles as representative samples to transform the intractable continuous MAP optimization into an efficient discrete combinatorial problem. Stein-MAP-Seq embeds SVGD into a Viterbi-style dynamic programming algorithm within a sequential variational inference framework that explicitly models the temporal dependencies inherent in dynamical systems. This novel combination leverages the mode-seeking yet repulsive behavior of SVGD to represent and separate multiple dominant modes, enabling efficient MAP sequence estimation via dynamic programming over the resulting discrete sample space. While the forward recursion incurs $\mathcal{O}(M^{2})$ computational complexity per time step, where $M$ denotes the number of particles, Stein-MAP-Seq achieves substantial efficiency gains by requiring far fewer particles ($M\leq 40$) compared to conventional particle filter-based MAP-Seq estimators that require hundreds to thousands of particles ($N\gg M$), resulting in dramatically lower computational cost in practice. Furthermore, since the particle updates in SVGD can be computed independently across samples, its inherent parallelism makes Stein-MAP-Seq naturally amenable to efficient implementation on modern parallel computing architectures, which are increasingly available in robotic platforms [34, 35, 36, 37, 38].
@@
Our approach fundamentally differs from traditional sampling-based methods by using SVGD’s deterministic particle evolution to generate high-quality samples that are then used as discrete states in a Viterbi-style dynamic programming algorithm. We rigorously evaluate the proposed method on challenging scenarios, including nonlinear dynamics with ambiguous measurements, pose estimation under unknown data associations with outliers, range-only localization under temporary unobservability, and high-dimensional manipulators with kinematic redundancy, demonstrating substantial improvements in estimation accuracy and robustness to multimodality over state-of-the-art baselines.
@@
The remainder of this paper is organized as follows: Section II formally defines the problem and presents our assumptions. Section III provides a brief overview of SVGD. Section IV details the proposed Stein-MAP-Seq algorithm. Section V presents estimation results across diverse scenarios and compares performance with existing methods. Finally, Section VI concludes the paper. Proofs and auxiliary results are provided in the appendices.
@@ II Problem Definition
Solving (2) is computationally intractable, particularly when $p(x_{0:T}|z_{1:T})$ exhibits multimodality and/or lacks a closed-form expression, making exact analytical computation impossible. To address this challenge, we propose a sequential variational inference framework to obtain an approximate MAP sequence solution. The details of the proposed approach are presented in Section IV. Before presenting those details, the next section provides background on the Stein Variational Gradient Descent method, which serves as a key component of our approach.
Input: The score function $\nabla_{x}\log p(x)$.
Goal: A set of particles $\{x_{i}\}_{i=1}^{N_{s}}$ that approximates $p(x)$.
For iteration l do
where $\hat{\phi}^{\star}(x_{i}^{(l)})$ is given in (III).
Algorithm 1 Stein Variational Gradient Descent
@@ III A Brief Overview of Stein Variational Gradient Descent method
Variational inference (VI) approximates a target distribution $p(x)$ using a simpler distribution $q^{\star}(x)$ found in a predefined set $\mathcal{Q}$ by minimizing the Kullback–Leibler (KL) [^1] KL divergence is defined as $\text{KL}[q(x)\parallel p(x)]=\int q(x)\log\frac{q(x)}{p(x)}dx$. divergence [39], i.e.,
@@
where the choice of the model space $\mathcal{Q}$ is critical for balancing accuracy and computational efficiency; however, achieving this balance remains challenging.
@@
The update $x_{i}^{(l+1)}=x_{i}^{(l)}+\epsilon\hat{\phi}^{\star}(x_{i}^{(l)})$ iteratively pushes particles toward $p(x)$, enabling expectation approximations via the empirical mean. Algorithm 1 summarizes the SVGD algorithm.
The samples generated by SVGD, denoted as $\{x_{i}\}_{i=1}^{N_{s}}$, serve a dual purpose in variational inference. First, they provide an empirical approximation of the optimal distribution $q^{\star}(x)$ in (3) as $q^{\star}(x)=\frac{1}{N_{s}}\sum\nolimits_{i=1}^{N_{s}}\delta(x-x^{i})$. Second, and equally important, these particles are commonly interpreted as representative samples from the target distribution $p(x)$ itself. Unlike traditional sampling methods such as MC or MCMC, which generate independent samples through stochastic processes, SVGD produces a deterministic set of interacting particles that collectively capture the structure of complex, potentially multimodal distributions through optimization-driven particle transport. This deterministic mechanism fundamentally differs from the random sampling variability inherent in Monte Carlo methods, offering computational advantages for downstream tasks, including efficient parallel computation and natural handling of multimodal posterior structures. Consequently, SVGD has emerged as a powerful sampling tool for approximate inference in high-dimensional and multimodal scenarios where traditional sampling methods face significant challenges.
Conventional SVGD addresses stationary inference problems, where the target distribution does not evolve over time. In contrast, state estimation in dynamical systems requires sequential inference over trajectories, where temporal dependencies between states must be preserved. In the next section, we extend SVGD to this sequential setting by developing a framework that leverages SVGD’s particle transport to approximate time-evolving conditional distributions while maintaining the Markov structure of dynamical systems, enabling efficient MAP sequence estimation via dynamic programming.
@@ IV Stein Maximum A Posteriori Sequence Estimation
In this section, we develop a MAP sequence estimator for the dynamical system (1) by integrating variational inference with dynamic programming. Computing the MAP trajectory $x_{0:T}^{\mathrm{MAP}}$ requires constructing an optimal distribution $q^{\star}(x_{0:T})$ whose support is concentrated on the (unknown) global maxima of the true posterior $p(x_{0:T}|z_{1:T})$, and subsequently identifying these modes. Formally, the problem can be stated as:
@@
Solving these optimization problems exactly presents significant challenges. When the dynamical system (1) is nonlinear, the resulting posterior $p(x_{0:T}|z_{1:T})$ becomes non-Gaussian and often multimodal, making closed-form representation impractical and exact analytical computation impossible. Moreover, the choice of $q(x_{0:T})$ significantly impacts estimation performance [16]. In the second optimization problem, finding the mode of $q^{\star}(x_{0:T})$ is intractable, since $q^{\star}(x_{0:T})$ typically resides on a high-dimensional continuous support.
To address these challenges, we propose Stein-MAP-Seq, a MAP-sequence estimator for dynamical systems (1) that preserves temporal dependencies via a sequential variational inference framework (Lemma IV.2). Stein-MAP-Seq integrates SVGD into a Viterbi-style dynamic programming algorithm, where SVGD’s deterministic gradient flow generates a compact discrete state representation, enabling efficient global MAP trajectory recovery through forward recursion and backtracking.
Our approach transforms the intractable continuous optimization in (5) into a tractable discrete problem through a two-stage strategy: first, SVGD generates particle-based approximations of the sequential posterior distributions, discretizing the continuous state space into a finite sample support; second, dynamic programming over this discrete particle support efficiently identifies the globally optimal MAP trajectory. This particle-based discretization fundamentally transforms the optimization from an intractable continuous search to an efficient discrete combinatorial problem.
The remainder of this section presents the theoretical foundations and algorithmic steps of our proposed Stein-MAP-Seq. We develop a series of marginalized conditional KL objectives (Lemma IV.2) that adapt SVGD from stationary inference to sequential trajectory estimation, preserving temporal dependencies through the Markov structure of dynamical systems. We then formalize the integration of these SVGD-generated particle approximations with Viterbi-style dynamic programming (Theorem IV.1), establishing the theoretical guarantees for MAP sequence recovery and analyzing computational complexity.